# Action Reasoning Models that can Reason in Space

### v2

@@ B.1 Implementation
To enable the model to learn to understand and output depth perception tokens, we follow the training scheme of LLaVA-AURORA (Bigverdi et al., 2025) by unfreezing the tokenizer embedding and lm head. For MolmoAct-7B-D, which uses Qwen2.5-7B, we simply replace the first 130 padding tokens with the depth perception tokens $\{\langle\mathrm{DEPTH\_START}\rangle,\langle\mathrm{DEPTH\_END}\rangle\}\;\cup\;\{\langle\mathrm{DEPTH}\_k\rangle\}_{k=1}^{128}$. However, for MolmoAct-7B-O, which uses Olmo2-7B, since it has less than 130 padding tokens, we first pad the tokenizer and lm head to its next multiple of 512, then replace the first 130 tokens with our depth perception tokens in the same way as MolmoAct-7B-D.
@@ B.2 GPU Cluster
[…]
@@
The servers are racked in Dynamic Density Cabinets333cirrascale.com/products-and-services/cabinet-technologies. Each cabinet houses five servers with dedicated cooling and power. Air circulates in a closed loop through an overhead plenum where it is cooled via heat transfer to water, enabling a datacenter PUE of 1.2. Under heavy utilization, H100 temperatures peak around $75^{\circ}\mathrm{C}$, with typical averages between $60^{\circ}\mathrm{C}$ and $65^{\circ}\mathrm{C}$.
@@
Pool Size ${\times}2 2×2$${\times}2$
@@
LR ViT ${\times}10^{-5} 1×1$${\times}10^{-5}$ $5{\times}10^{-6}$ $5{\times}10^{-6}$ $5{\times}10^{-4}$
LR Conn. ${\times}10^{-5} 1×1$${\times}10^{-5}$ $5{\times}10^{-6}$ $5{\times}10^{-6}$ $5{\times}10^{-4}$
LR LLM $2{\times}10^{-5}$ $2{\times}10^{-5}$ ${\times}10^{-5} 1×1$${\times}10^{-5}$ $5{\times}10^{-4}$

### v4

@@ B.1 Implementation
All of our collected data used for the mid- and post-training stages is recorded at 640480 px, which triggers the high-resolution cropping procedure described in Appendix A. By contrast, the OXE robot data used for pre-training has lower resolution, so no high-res crop is applied. To match OXE during pre-training, we downscale our collected images from 640480 to 320240 px while preserving the original aspect ratio. This alignment also reduces the number of vision tokens and accelerates training.
@@ B.2 GPU Cluster
Jupiter provides 1,024 NVIDIA H100 GPUs (80GB HBM3, 700W) across 128 servers. Each server has 2,,Intel Xeon Platinum8468 CPUs, 2TB DDR5 system memory, and 18 TB local NVMe storage.
@@
The servers are racked in Dynamic Density Cabinets333cirrascale.com/products-and-services/cabinet-technologies. Each cabinet houses five servers with dedicated cooling and power. Air circulates in a closed loop through an overhead plenum where it is cooled via heat transfer to water, enabling a datacenter PUE of 1.2. Under heavy utilization, H100 temperatures peak around $75\mathrm{C}$, with typical averages between $60\mathrm{C}$ and $65\mathrm{C}$.
@@
Pool Size ${\times}2 22$${\times}2$
@@
LR ViT ${\times}10^{-5} 11$${\times}10^{-5}$ $5{\times}10^{-6}$ $5{\times}10^{-6}$ $5{\times}10^{-4}$
LR Conn. ${\times}10^{-5} 11$${\times}10^{-5}$ $5{\times}10^{-6}$ $5{\times}10^{-6}$ $5{\times}10^{-4}$
LR LLM $2{\times}10^{-5}$ $2{\times}10^{-5}$ ${\times}10^{-5} 11$${\times}10^{-5}$ $5{\times}10^{-4}$